# A note on the scatteredness of reflection orders

### v1

@@ 3. Dense Subset
Let $(W,S)$ be an infinite, non-affine, finite rank, irreducible Coxeter system of rank greater than or equal to 3. Then $(W,S)$ has a reflection order which is not scattered.
@@
By Theorem 2.7.2 in [7], for an irreducible, infinite, non-affine Coxeter system $(W,S)$ of rank $\geq 3$, there exists a universal reflection subgroup $W^{\prime}$ of rank 3. Let the root basis for $W^{\prime}$ be $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$. This root basis must be linearly independent since otherwise the root system $W^{\prime}$ would be contained in a two dimensional space, and $W^{\prime}$ would be a dihedral group (rank 2 universal reflection subgroup). Let $V$ be the vector space in which the root system of $(W,S)$ resides. Choose a basis $B$ of $V$ that contains $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$. Extend the total order $\alpha_{1}\prec\alpha_{2}\prec\alpha_{3}$ to $B$. This induces a lexicographic reflection order $\prec_{\mathrm{reflex}}$ on $\Phi^{+}$ by Lemma 2.7. The restriction of $\prec_{\mathrm{reflex}}$ to $\Phi_{W^{\prime}}^{+}$ is precisely the reflection order discussed in Section 2.6, and by Theorem 3.5 it contains a subset to which the restriction of this reflection order is dense. Therefore the reflection order $\prec_{\mathrm{reflex}}$ is not scattered. ∎
@@
Let $(W,S)$ be an infinite irreducible Coxeter system. All reflection orders of $(W,S)$ are scattered if and only if $W$ is an affine Weyl group.
@@
By Theorem 4.4 from [16] (or by the main results in [3]), the order type of a reflection order on an affine Weyl group is a finite sum of copies of $\omega$ and $n]+\omega^{*}$. Such a total order is scattered by Proposition 2.17 from [14]. This together with Lemma 4.1 imply the assertion. ∎
We conclude this note by presenting two alternative, reflection order-theoretic characterizations (Theorem 4.3 and Theorem 4.6) that distinguish affine Weyl groups from all other Coxeter groups.
@@
Let $(W,S)$ be an infinite irreducible Coxeter system. The Coxeter group $W$ is an affine Weyl group if and only if it has a reflection order whose order type is $\omega+\omega^{*}.$
@@
Suppose that $W$ is an affine Weyl group. Let $\Phi_{0}^{+}$ be a positive system of the corresponding finite Weyl group $W_{0}.$ By the discussion preceding this theorem, the positive system $\Phi^{+}$ is the disjoint union of the “loop extension” $\widetilde{\Phi_{0}^{+}}$ of $\Phi_{0}^{+}$ and the “loop extension” $\widetilde{\Phi_{0}^{-}}$ of $\Phi_{0}^{-}$. By Theorem 3.12 of [15], $\widetilde{\Phi_{0}^{+}}$ and $\widetilde{\Phi_{0}^{-}}$ are both inversion sets of an infinite reduced word, i.e.
@@
(Here, an infinite reduced word is a sequence $s_{1}s_{2}\dots,s_{i}\in S$ such that any left prefix $s_{1}s_{2}\dots s_{k}$ of it is reduced.) Then one can easily see that the following total order is a reflection order:
@@
$\dots r_{1}r_{2}(\alpha_{r_{3}})\prec r_{1}(\alpha_{r_{2}})\prec\alpha_{r_{1}}$
which has order type $\omega+\omega^{*}.$
Conversely suppose that $(W,S)$ is not affine and assume to the contrary that it has a reflection order of the form
@@
Denote $I_{1}=\{\alpha_{1},\alpha_{2},\dots\}$ and $I_{2}=\{\beta_{1},\beta_{2}.\dots\}.$ By Theorem 2.7.2 in [7], $W$ has a universal reflection subgroup $W^{\prime}$ of rank 3. Denote the simple roots of $W^{\prime}$ by $\gamma_{1},\gamma_{2},\gamma_{3}$. By pigeonhole’s principle, at least two of them have to be contained in the same $I_{i}$. Without loss of generality, assume $\gamma_{1}$ and $\gamma_{2}$ are contained in $I_{1}$. Then the roots of the parabolic subgroup $U$ (of $W^{\prime}$) generated by $s_{\gamma_{1}}$ and $s_{\gamma_{2}}$ must all be contained in $I_{1}$. Furthermore the restriction of $\prec$ to $U$ has the order type $\omega+\omega^{*}$. But $I_{1}$ itself has the order type $\omega$. A contradiction. ∎
The following proposition shows that the order type of a reflection order of a finite-rank Coxeter group always begins with $\omega$ and ends with $\omega^{*}$. This implies that a reflection order is never dense itself.
@@
Let $\prec$ be a reflection order of $W$. Write $\Pi=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n}\}$. The simple reflection corresponding to $\alpha_{i}$ is written as $s_{i}$. Assume that restricted to $\Pi$, one has
@@
We show that $\alpha_{1}$ must be the minimum element with respect to $\prec$. Take a root $\beta=s_{i_{k}}s_{i_{k-1}}\dots s_{i_{1}}(\alpha_{j})\neq\alpha_{1}.$ Note that $\alpha_{1}\preceq\alpha_{j}$. Now we assume that $\alpha_{1}\preceq s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j})$. Since $s_{i_{t+1}}s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j})=s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j})+k\alpha_{i_{t+1}}$ and $\alpha_{1}\preceq\alpha_{i_{t+1}}$, we conclude that $\alpha_{1}\preceq s_{i_{t+1}}s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j}).$ So by induction $\alpha_{1}\prec\beta$.
Now we construct upper $s_{1}$-conjugate of $\prec$. One obtains another reflection order with minimum $s_{j_{1}}$. Then we construct $s_{j_{1}}$-conjugate to the obtained reflection order. One obtains another reflection order with minimum $s_{j_{2}}$. Repeatedly carry out this process and we can now conclude that the reflection order $\prec$ starts with
@@
Therefore one sees that $\prec$ has an initial interval of order type $\omega$. Since the opposite order of a reflection order is clearly still a reflection order, $\prec$ has a final interval of order type $\omega^{*}.$
The immediate corollary of the above Proposition is
@@
Let $(W,S)$ be a finite rank infinite Coxeter system. Then every reflection order of $(W,S)$ has an initial interval that is the inversion set of an infinite reduced word.
@@
Let $(W,S)$ be an infinite irreducible Coxeter system. Then the Coxeter group $W$ is an affine Weyl group if and only if, for every dihedral reflection subgroup $U$ and any reflection order $\prec$ of $(W,S)$, there are only finitely many pairs of positive roots $\gamma_{1},\gamma_{2}\in\Phi_{U}^{+}$ such that $\gamma_{1}$ is the immediate predecessor of $\gamma_{2}$ (under the restriction of $\prec$ to $\Phi_{U}^{+}$) and the interval $\gamma_{1},\gamma_{2}]$ under $\prec$ is infinite.
@@
First assume that $(W,S)$ is affine. The proof in this case proceeds along the same lines as the proof of Lemma 3.3 in [16]. The restriction of any reflection order on a dihedral reflection subgroup $U$ of $W$ takes the form
@@
Note that $\Phi_{0}$ is finite. If $\alpha_{i},\alpha_{i+1}]$ (resp. $\beta_{i+1},\beta_{i}]$) is infinite, then its intersection with $\widetilde{\eta}$ must be infinite for some $\eta\in\Phi_{0}$. Furthermore since the restriction of any reflection order on $\widetilde{\{\pm\eta\}}$ (which is the positive system of a dihedral reflection subgroup generated by $s_{\eta_{0}}$ and $s_{(-\eta)_{0}}$) is again a reflection order on $\widetilde{\{\pm\eta\}}$, the intersection of $\alpha_{i},\alpha_{i+1}]$ (resp. $\beta_{i+1},\beta_{i}]$) with $\widetilde{\eta}$ has to be $\{\eta+k\delta\mid\,k\geq k_{0}\}$ for some integer $k_{0}$. Again by the finiteness of $\Phi_{0}$, there can only be finite many such infinite intervals $\alpha_{i},\alpha_{i+1}]$ (resp. $\beta_{i+1},\beta_{i}]$).
Conversely assume that $(W,S)$ is not affine. By Theorem 2.7.2 in [7], it has a universal reflection subgroup $W^{\prime}$ of rank 3. Similar to the proof of Lemma 4.1, There exists a reflection order $\prec$ of $W$, when restricted to (the set of the positive roots of) $W^{\prime}$, coincides with the lexicographic reflection order described in Section 3. By Lemma 3.6, there exists a dihedral reflection subgroup $U$ of $W^{\prime}$ for which, with respect to the reflection order $\prec$, the interval $\gamma_{1},\gamma_{2}]$ is infinite for every pair of positive roots $\gamma_{1},\gamma_{2}$ such that $\gamma_{1}$ is the immediate predecessor of $\gamma_{2}$. ∎
@@ Acknowledgment
The first author is grateful to Professor Matthew Dyer for very helpful communication and, in particular, for suggesting the idea that led to the proof of Lemma 3.2.

### v2

@@ 4. Non-scatteredness of reflection orders of non-affine, infinite Coxeter groups
Let $(W,S)$ be an infinite, non-affine, finite rank, irreducible Coxeter system of rank greater than or equal to 3. Then $(W,S)$ has a reflection order that is not scattered.
@@
By Theorem 2.7.2 in [6], if $(W,S)$ is an irreducible, infinite, non-affine Coxeter system of rank $\geq 3$, then $W$ has a universal reflection subgroup $W^{\prime}$ of rank 3. Let the root basis for $W^{\prime}$ be $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$. This root basis must be linearly independent since otherwise the root system $W^{\prime}$ would be contained in a two dimensional space, and $W^{\prime}$ would be a dihedral group (i.e. a rank 2 universal reflection subgroup). Let $V$ be the vector space in which the root system of $(W,S)$ resides. Choose a basis $B$ of $V$ containing $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$. Extend the total order $\alpha_{1}\prec\alpha_{2}\prec\alpha_{3}$ to a total order on $B$, which further induces a lexicographic reflection order $\prec_{\mathrm{reflex}}$ on $\Phi^{+}$ by Lemma 2.7. The restriction of $\prec_{\mathrm{reflex}}$ to $\Phi_{W^{\prime}}^{+}$ is precisely the reflection order discussed in Section 2.6. By Theorem 3.5, $\Phi_{W^{\prime}}^{+}$ has a subset to which the restriction of this reflection order is dense. Therefore, the reflection order $\prec_{\mathrm{reflex}}$ is not scattered. ∎
@@
Let $(W,S)$ be an infinite, irreducible Coxeter system. Any reflection order of $(W,S)$ is scattered if and only if $W$ is an affine Weyl group.
@@
By Theorem 4.4 in [15] (or by the main results in [2]), the order type of a reflection order on an affine Weyl group is a finite sum of copies of $\omega$ and $n]+\omega^{*}$. Such a total order is scattered by Proposition 2.17 in [13]. Together with Lemma 4.1, this implies the assertion. ∎
Next we present two alternative, reflection order-theoretic characterizations (Theorem 4.3 and Theorem 4.7) that distinguish affine Weyl groups from all other Coxeter groups.
@@
Let $(W,S)$ be an infinite, irreducible Coxeter system. The Coxeter group $W$ is an affine Weyl group if and only if it has a reflection order with order type $\omega+\omega^{*}.$
@@
Suppose that $W$ is an affine Weyl group. Let $\Phi_{0}^{+}$ be a positive system of the corresponding finite Weyl group $W_{0}.$ By the preceding discussion, the positive system $\Phi^{+}$ is the disjoint union of the “loop extension” $\widetilde{\Phi_{0}^{+}}$ of $\Phi_{0}^{+}$ and the “loop extension” $\widetilde{\Phi_{0}^{-}}$ of $\Phi_{0}^{-}$. By Theorem 3.12 in [14], both $\widetilde{\Phi_{0}^{+}}$ and $\widetilde{\Phi_{0}^{-}}$ are inversion sets of an infinite reduced word, i.e.
@@
Here, an infinite reduced word is a sequence $s_{1}s_{2}\dots,s_{i}\in S$ such that any left prefix $s_{1}s_{2}\dots s_{k}$ of it is reduced. It is then easy to see that the following total order is a reflection order:
@@
$\dots r_{1}r_{2}(\alpha_{r_{3}})\prec r_{1}(\alpha_{r_{2}})\prec\alpha_{r_{1}},$
which has the order type $\omega+\omega^{*}.$
Conversely, suppose that $(W,S)$ is not affine and assume to the contrary that it has a reflection order of the form
@@
Write $I_{1}=\{\alpha_{1},\alpha_{2},\dots\}$ and $I_{2}=\{\beta_{1},\beta_{2}.\dots\}.$ By Theorem 2.7.2 in [6], $W$ has a universal reflection subgroup $W^{\prime}$ of rank 3. Denote the simple roots of $W^{\prime}$ by $\gamma_{1},\gamma_{2},\gamma_{3}$. By the pigeonhole principle, at least two of these have to be contained in the same $I_{i}$. Without loss of generality, assume that $\gamma_{1}$ and $\gamma_{2}$ are contained in $I_{1}$. In this case, the roots of the parabolic subgroup $U$ (of $W^{\prime}$) generated by $s_{\gamma_{1}}$ and $s_{\gamma_{2}}$ must all be contained in $I_{1}$. Furthermore the restriction of $\prec$ to $U$ has the order type $\omega+\omega^{*}$. However, $I_{1}$ itself has the order type $\omega$. A contradiction. ∎
The following proposition shows that the order type of a reflection order of a finite-rank Coxeter group always begins with $\omega$ and ends with $\omega^{*}$. This implies that although a reflection order may contain a dense subset, it can never itself be dense.
@@
Let $\prec$ be a reflection order of $W$. Write the root basis $\Pi=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n}\}$. The simple reflection corresponding to $\alpha_{i}$ is written as $s_{i}$. Assume that restricted to $\Pi$, one has
@@
We show that $\alpha_{1}$ must be the minimum element with respect to $\prec$. Take a root $\beta=s_{i_{k}}s_{i_{k-1}}\dots s_{i_{1}}(\alpha_{j})\neq\alpha_{1}.$ Note that $\alpha_{1}\preceq\alpha_{j}$. Now we assume that $\alpha_{1}\preceq s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j})$. Since $s_{i_{t+1}}s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j})=s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j})+k\alpha_{i_{t+1}}$ and $\alpha_{1}\preceq\alpha_{i_{t+1}}$, we conclude that $\alpha_{1}\preceq s_{i_{t+1}}s_{i_{t}}s_{i_{t-1}}\dots s_{i_{1}}(\alpha_{j}).$ Therefore, by induction $\alpha_{1}\prec\beta$.
We now construct the upper $s_{1}$-conjugate of $\prec$. This gives another reflection order with minimum $s_{j_{1}}$. Then we construct the upper $s_{j_{1}}$-conjugate of the obtained reflection order. This yields another reflection order with minimum $s_{j_{2}}$. By repeatedly carrying out this process, we can now conclude that the reflection order $\prec$ starts with
@@
Therefore, it can be seen that $\prec$ has an initial interval of the order type $\omega$. Since the backward order of a reflection order is still a reflection order, $\prec$ also has a final interval of order type $\omega^{*}.$
The immediate consequence of the above Proposition is the following Corollary.
@@
Let $(W,S)$ be a finite rank, infinite Coxeter system. Every reflection order of $(W,S)$ has an initial interval (resp. final interval) that is the inversion set of an infinite reduced word.
Remark 4.6.
Proposition 4.4 and Corollary 4.5 no longer hold when the rank of the Coxeter system is infinite. Let $S=\{s_{1},s_{2},s_{3},\dots\}$ and
$W=\langle S\mid\,s_{i}^{2}=e,(s_{i}s_{i+1})^{3}=e,\text{for all $i=1,2,\dots$},(s_{i}s_{j})^{2}=e,\text{for $|i-j|\geq 2$}\rangle.$
Then $(W,S)$ is of type $A_{\infty}$ and $W$ is the group of permutations with finite support of the natural numbers. Write $\alpha_{i}=\alpha_{s_{i}}$. The following total order $\prec$ is a reflection order:
$\alpha_{1}\prec\alpha_{1}+\alpha_{2}\prec\alpha_{2}\prec\alpha_{1}+\alpha_{2}+\alpha_{3}\prec\alpha_{2}+\alpha_{3}\prec\alpha_{3}\prec\dots$
$\alpha_{1}+\alpha_{2}+\dots+\alpha_{n}\prec\alpha_{2}+\alpha_{3}+\dots+\alpha_{n}\prec\dots\prec\alpha_{n-1}+\alpha_{n}\prec\alpha_{n}\prec\dots$
This reflection order has the order type of $\omega$. The opposite order $\prec^{*}$ is also a reflection order with the order type $\omega^{*}$. Neither of these two phenomena occurs in finite rank. Applying upper and lower $s$-conjugates successively to $\prec$, one obtains reflection orders with the order types $\omega+[n]$ and $n]+\omega^{*}$.
On the other hand, the order type of a reflection order of $(W,S)$ can still be $\omega+\omega^{*}$. Let
$w_{n}=\begin{cases}s_{n}s_{n-1}\dots s_{1},&\text{if $n$ is even},\\ s_{n}s_{n-1}\dots s_{2},&\text{if $n$ is odd}.\end{cases}$
The concatenation $u=w_{2}w_{3}w_{4}w_{5}\dots$ is an infinite reduced word and let $u=r_{1}r_{2}r_{3}\dots$ one of its reduced expression. Write
$\beta_{i}=r_{1}r_{2}\dots r_{i-1}(\alpha_{r_{i}}).$
Then the total order
$\alpha_{1}\prec\alpha_{3}\prec\alpha_{5}\prec\dots\beta_{3}\prec\beta_{2}\prec\beta_{1}$
is a reflection order with the order type $\omega+\omega^{*}.$ Indeed one can even assign arbitrary total order to the initial interval $\{\alpha_{1},\alpha_{3},\alpha_{5},\dots\}$. Therefore the order type $\eta+\omega^{*}$ can be the order type of a reflection order of $(W,S)$. Therefore we can conclude that for $(W,S)$, not every reflection order is scattered.
@@
Let $(W,S)$ be an infinite irreducible Coxeter system. Then the Coxeter group $W$ is an affine Weyl group if and only if, for every dihedral reflection subgroup $U$ and any reflection order $\prec$ of $(W,S)$, there are only finitely many pairs of positive roots $\gamma_{1},\gamma_{2}\in\Phi_{U}^{+}$ such that, under the restriction of $\prec$ to $\Phi_{U}^{+}$, $\gamma_{1}$ is the immediate predecessor of $\gamma_{2}$ and the interval $\gamma_{1},\gamma_{2}]$ under $\prec$ is infinite.
@@
First, assume that $(W,S)$ is affine. The proof in this case proceeds along the same lines as that of Lemma 3.3 in [15]. The restriction of any reflection order to a dihedral reflection subgroup $U$ of $W$ takes the form
@@
Note that $\Phi_{0}$ is finite. If $\alpha_{i},\alpha_{i+1}]$ (resp. $\beta_{i+1},\beta_{i}]$) is infinite, then its intersection with $\widetilde{\eta}$ must be infinite for some $\eta\in\Phi_{0}$ because of the finiteness of $\Phi_{0}$. Furthermore since the restriction of any reflection order on $\widetilde{\{\pm\eta\}}$ (which is the positive system of a dihedral reflection subgroup generated by $s_{\eta_{0}}$ and $s_{(-\eta)_{0}}$) is again a reflection order on $\widetilde{\{\pm\eta\}}$, the intersection of $\alpha_{i},\alpha_{i+1}]$ (resp. $\beta_{i+1},\beta_{i}]$) with $\widetilde{\eta}$ has to be $\{\eta+k\delta\mid\,k\geq k_{0}\}$ for some integer $k_{0}$. Again by the finiteness of $\Phi_{0}$, there can be only a finite number of such infinite intervals $\alpha_{i},\alpha_{i+1}]$ (resp. $\beta_{i+1},\beta_{i}]$).
Conversely, assume that $(W,S)$ is not affine. By Theorem 2.7.2 in [6], it has a universal reflection subgroup $W^{\prime}$ of rank 3. Similar to the proof of Lemma 4.1, there exists a reflection order $\prec$ of $W$ that, when restricted to (the set of the positive roots of) $W^{\prime}$, coincides with the lexicographic reflection order described in Section 3. By Lemma 3.6, there exists a dihedral reflection subgroup $U$ of $W^{\prime}$ such that, with respect to the reflection order $\prec$, the interval $\gamma_{1},\gamma_{2}]$ is infinite for every pair of positive roots $\gamma_{1},\gamma_{2}\in\Phi_{U}^{+}$ for which $\gamma_{1}$ is the immediate predecessor of $\gamma_{2}$ under the restriction of $\prec$ to $\Phi_{U}^{+}$. ∎
We conclude the note by presenting a reflection-order-theoretic characterization of non-affine irreducible Coxeter systems among all Coxeter systems.
Theorem 4.8.
An infinite, irreducible Coxeter system $(W,S)$ is non-affine if and only if, for any reflection order $\prec$ on $\Phi_{W}^{+}$ and any positive integer $k$, there exists a subset $E\subset\Phi_{W}^{+}$ such that the restriction of $\prec$ to $E$ has the order type $(\omega+\omega^{*})[k]$.
Proof.
Suppose that $(W,S)$ is a non-affine, infinite, irreducible Coxeter system. As in the previous proofs, $W$ contains a universal reflection subgroup $W^{\prime}$ of rank $3$ by Theorem 2.7.2 of [6]. Any reflection order $\prec$ on $W$ restricts to a reflection order on $W^{\prime}$. The existence of the subset $E$ therefore follows from Lemma 3.8.
Conversely, suppose that $(W,S)$ is irreducible and affine of rank $n$ (that is, the associated finite Weyl group has Coxeter rank $n$). By Theorem 4.4 of [15], the order type of any reflection order of $W$ is a sum of copies of $\omega$ and $t]+\omega^{*}$, and the resulting sequence of these terms forms a subsequence of a $2n$-extended Dyck word. Consequently, the order type of $\prec$ contains at most $n$ copies of $\omega$ and at most $n$ copies of $\omega^{*}$. This completes the proof. ∎
@@ Acknowledgment
The first author is grateful to Professor Matthew Dyer for very helpful communication and, particularly, for suggesting the idea that led to the proof of Lemma 3.2.